# Non-linear conductances of Galton–Watson trees and application to the (near) critical random cluster model

### v1

@@ 1.3 Main results: $p$-conductances of Galton–Watson trees
From now on, we fix $p>1$ so we will suppress it from notations and write $C_{n}(u)$ instead of $C_{n}^{(p)}(u)$. We wish to estimate the $p$-conductance of the Galton–Watson tree $T_{n}$, that is the $p$-conductance between the root and the leaves of $T_{n}$:
$C_{n}:=C_{n}(\rho)=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})\,,$
which is a random variable since it depend on the realization of the tree $T$. In the following, we focus on the case where the tree is equipped with resistances $(R_{v})_{v\in T_{n}}$ that are of the form:
@@
where $|v|$ is the distance between $v$ and the root, and $R=R_{n}>0$ is a number which is fixed along $T_{n}$, but that may depend on $n$. The choice (1.7) may seem rather restrictive but is the one that appears naturally in the statistical mechanics models that we are interested in; then $R_{n}$ is related to the inverse temperature of the model (the fact that $R_{n}$ may depend on $n$ will allow us to study the critical window for the random cluster model, as will become clear in Section 2). It is also closely related to the choice of conductances considered in [9] (or [1]), where authors estimate the standard $2$-conductance (or $2$-resistance) with resistances $R_{v}=\xi_{v}\,m^{|v|}$, for i.i.d. random variables $(\xi_{v})_{v\in T}$ — here we focus on the case where $\xi_{v}\equiv 1$ but we allow more generality in the growth rate of resistances in view of the applications we have in mind, see Section 2.
More generally, we consider a concave function $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ verifying (1.5) and the recursion (1.3) with resistances $(R_{v})_{v\in T_{n}}$ defined as in (1.7): with this choice of resistances, we have for all $u\in T_{n}\setminus\partial T_{n}$
@@
with boundary condition $B_{n}(u)\equiv b_{0}$ for $u\in\partial T_{n}$. We also denote $B_{n}:=B_{n}(\rho)$; in particular, we have $B_{n}=C_{n}$ if $g(x)=x/(1+x^{s})^{1/s}$ and $b_{0}=1$.
Let us stress once more that the resistances are $R_{v}=(R_{n})^{-|v|}$, where $R_{n}$ is a parameter that is fixed along the tree $T_{n}$ and that might depend on $n$. In particular, except when $R_{n}\equiv R$, the recursion (1.8) that defines $B_{n}$ is different from the one that defines $B_{n+1}$. In fact, one may think of $B_{n}$ as belonging to a (tree-)triangular array of recursions $((B_{n}(u))_{u\in T_{n}},n\geqslant 0)$.
@@
To anticipate a bit, let us notice that we have $a_{n}=f_{n}(mR_{n})$, with $f_{n}(x)=x^{-s}\frac{x^{-ns}-1}{x^{-s}-1}$. In particular, we easily get that
[…]
where we have used the notation $a_{n}\asymp a_{n}^{\prime}$ if the ratio $a_{n}/a_{n}^{\prime}$ is bounded from above and from below by two universal constants. Note that the first line also includes the case where $R_{n}\to 0$, but we will mostly focus on the case where $\liminf_{n\to\infty}R_{n}>0$. We refer to the regime $mR_{n}\equiv 1$ as the critical case and one can interpret $|mR_{n}-1|=O(\frac{1}{n})$ as being in the critical window.
We obtain different results depending on whether $Z\sim\mu$ admits or not a finite moment of order $q$; similarly to what happens for the critical behavior of the Ising model on random trees, see [17]. We start with the case where $\mathbf{E}[Z^{q}]<+\infty$, where our results are much sharper; we then turn to the case where $\mathbf{E}[Z^{q}]=+\infty$.
1.3.1 Case of a finite moment of order $q$
Our first result gives a general estimate on $\mathbf{E}[B_{n}]$; in particular, thanks to Proposition 1.4, we may focus on the $p$-conductance $C_{n}$ of the tree $T_{n}$. We first give a result in the case where the offspring distribution admits a finite moment of order $q$.
@@ 1.3.1 Case of a finite moment of order $q$
Let $p>1$, let $q=\frac{p}{p-1}$ be its conjugate exponent and let $s=q-1=\frac{1}{p-1}$. Assume that $Z\sim\mu$ admits a finite moment of order $q$. Then, there is a constant $c_{p}$ (that depend only on the law $\mu$ and on $p$) such that
@@
Let us note that when $Z$ only admits a moment of order $r<q$, then Lemma 1.5 combined with Theorem 1.6 gives the bound
@@
[…]
@@
Let $p>1$, let $q=\frac{p}{p-1}$ be its conjugate exponent and $s=q-1$. Assume that $Z\sim\mu$ admits a finite moment of order $q$. Then, if $\limsup_{n\to\infty}mR_{n}\leqslant 1$ and $\inf_{n\geqslant 0}R_{n}>0$, we have, as $n\to\infty$
$\frac{B_{n}}{\mathbf{E}[B_{n}]}\xrightarrow[n\to\infty]{L^{p}}W\,,$
where $W:=\lim_{n\to\infty}\frac{1}{m^{n}}Z_{n}$ is the a.s. limit of the usual martingale associated to the branching process. If $R_{n}\equiv R$ with $R\in(0,m^{-1}]$, then the convergence also holds almost surely.
@@
Let $p>1$, let $q=\frac{p}{p-1}$ be its conjugate exponent and let $s=q-1$. Assume that $Z\sim\mu$ admits a finite moment of order $q$ and that the following limit exists $\kappa_{g}:=\lim_{x\downarrow 0}\frac{1}{x^{q}}(x-g(x))$ in $(0,+\infty)$. Then, if $\lim_{n\to\infty}mR_{n}=\vartheta\in(0,1]$, we have, as $n\to\infty$,
$\mathbf{E}[B_{n}]\sim\alpha_{p}(\vartheta)(a_{n})^{-1/s}\,,$
for some constant $\alpha_{p}(\vartheta)$, with $\alpha_{p}(1):=(s\kappa_{g}\mathbf{E}[W^{q}])^{-1/s}$ in the case where $\lim_{n\to\infty}mR_{n}=1$.
Let us mention that Proposition 1.9 improves in particular the result [9, Thm. 1.2] which considers some additional source of randomness on the resistances, but treats only the case of linear ($p=q=2$) conductances in the critical case $mR_{n}\equiv 1$ and requires a moment of order $3$ (i.e. $q+1$) for the branching process.
1.3.2 Case of an infinite moment of order $q$
In the case of an infinite moment of order $q$, we need assume some (either upper or lower) bounds on the truncated $q$-moment of $Z\sim\mu$. More precisely, we assume that there is some $\alpha\in(1,q]$ and some slowly varying function $L(\cdot)$ such that, for $x\geqslant 1$
$\mathbf{E}[(Z\wedge x)^{q}]\geqslant c_{1}L(x)x^{q-\alpha}$ (1.12)
$\mathbf{E}[(Z\wedge x)^{q}]\leqslant c_{2}L(x)x^{q-\alpha}$ (1.13)
@@ 1.3.2 Case of an infinite moment of order $q$
Let $h$ be a increasing function which verifies $h(x)\sim L(1/x)x^{\alpha-1}$ as $x\downarrow 0$. We denote by $h^{-1}$ its inverse and let us stress that $h^{-1}(x)\sim\tilde{L}(x)x^{1/(\alpha-1)}$ as $x\downarrow 0$, for some slowly varying function $\tilde{L}$, see [8, §1.5.7]. Then, define
[…]
to be compared with (1.11), which corresponds to having $h(x)=x^{q-1}$. Let us also define $\tilde{\gamma}_{n}$ by setting $\tilde{\gamma}_{n}=\gamma_{n}$ if $mR_{n}\geqslant 1-\frac{1}{n}$ and $\tilde{\gamma}_{n}:=h^{-1}(1/n)(n(1-mR_{n}))^{1/s}(mR_{n})^{n}$ if $mR_{n}\leqslant 1-\frac{1}{n}$.
In practice, in the case where $R_{n}\equiv R$ and if $L(x)\equiv 1$ in (1.12)-(1.13), then we have: (i) $\gamma_{n}=\tilde{\gamma}_{n}\asymp(mR-1)^{1/(\alpha-1)}$ if $mR>1$; (ii) $\gamma_{n}=\tilde{\gamma}_{n}\asymp n^{-1/(\alpha-1)}$ if $mR=1$; (iii) $\gamma_{n}\asymp(mR)^{n}$ and $\tilde{\gamma}_{n}\asymp n^{\frac{1}{q-1}-\frac{1}{\alpha-1}}\gamma_{n}$ if $mR<1$.
@@
Assume that $0<\inf_{n}mR_{n}\leqslant\sup_{n}mR_{n}\leqslant 2$ and let $\gamma_{n}$ be as in (1.14). Then there are constants $c,c^{\prime}$ such that:
@@
Let us also show that the conditions (1.12)-(1.13) are weaker than the tail conditions (1.15)-(1.16). Indeed, on one hand, we have $\mathbf{E}[(Z\wedge x)^{q}]\geqslant x^{q}\mathbf{P}(Z>x)$, so (1.13) implies the upper bound $\mathbf{P}(Z>x)\leqslant c_{2}L(x)x^{-\alpha}$, but this is not optimal in the case $\alpha=q$. However, in the case $\alpha<q$, (1.13) is equivalent to (1.16) with $\hat{L}=L$. On the other hand, to obtain a lower bound on $\mathbf{P}(Z>x)$, one needs to have $\alpha<q$ and both (1.12)-(1.13); simply write $\mathbf{E}[(Z\wedge Ax)^{q}]\leqslant(Ax)^{q}\mathbf{P}(Z>x)+\mathbf{E}[(Z\wedge x)^{q}]$.
@@ 1.4 Organisation of the paper
In section 2 we introduce the main motivation for our results on non-linear conductances: the random cluster model (RCM) on Galton–Watson trees. In 2.2, we present the tree recursion arising in the model (we postpone the proof to Appendix A) and, in Section 2.3, we apply our main results on $p$-conductances in order to detail the critical behaviour of the RCM on a quenched GW tree.
@@
  • In Section 6, we prove 1.10, which treats the moments of $C_{n}$ in the case $Z\sim\mu$ has an infinite moment of order $p$.
  • In Appendix A we prove the RCM tree recursion presented in Section 2.2.
  • In Appendix B we provide some technical proofs regarding branching processed with heavy tails, which are necessary for the estimates of Theorem 1.10.
2 The (critical) random cluster model on a quenched GW tree
The Random Cluster Model (RCM) or FK-percolation, introduced by Fortuin and Kasteleyn in [19], unifies a number of models in statistical physics, including the percolation model, the Ising or $\mathtt{q}$-states Potts models and the uniform spanning tree. Some of these models, when considered on trees, exhibit recursions as in (1.3), in particular the connection (sometimes called survival) probability of the RCM on a tree, see (2.2). We give a quick presentation of the model below, but we refer to [21] for a more complete overview of the RCM and its connection to other models; see also [18, Ch. 1].
Let us mention that one of the motivation for considering statistical mechanics models on trees is to apply the results to general random tree-like graphs, see e.g. [35] for an overview (see also [13, 16] in the context of the Ising model).

### v2

@@ 1.3 Main results: $p$-conductances of Galton–Watson trees
From this point on, we will work with a fixed $p>1$ so we omit it from the notation, writing $C_{n}(u)$ in place of $C_{n}^{(p)}(u)$. We wish to estimate the $p$-conductance of the Galton–Watson tree $T_{n}$, that is, the $p$-conductance between the root and the leaves of $T_{n}$:
$C_{n}:=C_{n}(\rho)=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})\,.$
Note that this is a random variable since it depend on the realization of the tree $T_{n}$. In the following, we focus on the case where the tree is equipped with resistances $(R_{v})_{v\in T_{n}}$ that are of the form:
@@
where $|v|$ is the distance between $v$ and the root, and $R=R_{n}>0$ is a number which is fixed along $T_{n}$, but that may depend on $n$. The choice (1.7) may seem rather restrictive but this is what appears naturally in the statistical mechanics models that we are interested in; then $R_{n}$ is related to the inverse temperature of the model (the fact that $R_{n}$ may depend on $n$ will allow us to study the critical window for the random cluster model, as will become clear in Section 2). It is also closely related to the choice of conductances considered in [1, 10], where the authors estimate the standard $2$-conductance (or $2$-resistance) with resistances $R_{v}=\xi_{v}\,m^{|v|}$, for i.i.d. random variables $(\xi_{v})_{v\in T}$ — here we focus on the case where $\xi_{v}\equiv 1$ but we allow more generality in the growth rate of resistances in view of the applications we have in mind, see Section 2.
More generally, we will consider recursions (1.4) with a function $f_{v}\equiv R_{n}g$ that does not depend on $v$, for some function $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ that is concave and bounded, with $g^{\prime}(0+)=1$ — the latter is simply a normalization choice. More precisely, we study recursions of the type
@@
with boundary condition $B_{n}(u)\equiv+\infty$ for $u\in\partial T_{n}$. We also denote $B_{n}:=B_{n}(\rho)$. In particular, in view of the recursion (1.3), we have that $B_{n}(u)=C_{n}(u)$} if we take the function $g(x)=x(1+x^{s})^{-1/s}$. Our main assumption is that $g$ is concave and bounded as mentioned, and verifies a bound similar to (1.5): there is some $s>0$ and some constants $\kappa_{1}$, $\kappa_{2}$ such that for all $x>0$
$\frac{x}{(1+\kappa_{1}x^{s})^{1/s}}\leqslant g(x)\leqslant\frac{x}{(1+\kappa_{2}x^{s})^{1/s}}\,.$ (1.9)
This is for instance verified in the random cluster model with cluster weight $\mathtt{q}$ (with $s=2$ if $\mathtt{q}\in(0,2)$} and $s=3$ for $\mathtt{q}=2$), see Section 2 below.
We emphasize once again that the resistances are given by $R_{v}=(R_{n})^{-|v|}$, where $R_{n}$ is a parameter that remains fixed throughout the tree $T_{n}$, but may depend on $n$. In particular, except when $R_{n}\equiv R$, the recursion (1.8) that defines $B_{n}$ is different from the one that defines $B_{n+1}$. In fact, one may think of $B_{n}$ as belonging to a (tree-)triangular array of recursions $((B_{n}(u))_{u\in T_{n}},n\geqslant 0)$.
@@
To anticipate a bit, let us notice that we have $a_{n}=f_{n}(mR_{n})$, with $f_{n}(x)=x^{-s}\frac{x^{-ns}-1}{x^{-s}-1}$. In particular, we easily get that [^2] We used the standard notation $a_{n}\asymp a_{n}^{\prime}$ if the ratio $a_{n}/a_{n}^{\prime}$ is bounded from above and from below by two universal constants.
 $\begin{array}[]{lll}\text{ if }mR_{n}\in\big(0,1-\tfrac{1}{n}\big]&\qquad a_{n}\asymp(1-(mR_{n})^{s})^{-1}(mR_{n})^{-ns}\,,\\[2.0pt] \text{ if }mR_{n}\in\big[1-\tfrac{1}{n},1+\tfrac{1}{n}\big]&\qquad a_{n}\asymp n\,,\\[2.0pt] \text{ if }mR_{n}\in\big[1+\tfrac{1}{n},2]&\qquad a_{n}\asymp((mR_{n})^{s}-1)^{-1}\,.\end{array}$ (1.11)
Note that the first line also includes the case where $R_{n}\to 0$, but we will mostly focus on the case where $\liminf_{n\to\infty}R_{n}>0$. We refer to the regime $mR_{n}\equiv 1$ as the critical case and one can interpret $|mR_{n}-1|=O(\frac{1}{n})$ as the critical window.
We obtain different results depending on whether $Z\sim\mu$ admits or not a finite moment of order $q:=s+1$; similarly to what happens for the critical behavior of the Ising model on random trees, see [19]. We start with the case where $\mathbf{E}[Z^{q}]<+\infty$, where our results are much sharper; we then turn to the case where $\mathbf{E}[Z^{q}]=+\infty$.
1.3.1 Case of a finite moment of order $q=s+1$
Our first result gives a general estimate on $\mathbf{E}[B_{n}]$; in particular, thanks to Proposition 1.4, we may focus on the $p$-conductance $C_{n}$ of the tree $T_{n}$, with $p=\frac{s+1}{s}$. We first present a result in the case where the offspring distribution has a finite moment of order $q:=\frac{p}{p-1}=s+1$ (note that $q$ is the conjugate exponent of $p$), and subsequently derive bounds under the weaker assumption of a finite moment of order $r<q$.
@@ 1.3.1 Case of a finite moment of order $q=s+1$
Let $p>1$ and let $s=\frac{1}{p-1}$. Assume that $Z\sim\mu$ admits a finite moment of order $q=\frac{p}{p-1}=s+1$. Then, there is a constant $c_{p}$ (that depend only on the law $\mu$ and on $p$) such that
@@
Let us note that when $Z$ only admits a moment of order $r<q$, then Lemma 1.5 (note that the monotonicity is reversed when considering the conjugate exponents) combined with Theorem 1.6 gives the bound
@@
 $\begin{array}[]{lll}\text{ if }mR_{n}\in\big(0,1-\tfrac{1}{n}\big]&\qquad\mathbf{E}[B_{n}]\asymp(1-(mR_{n})^{s})^{1/s}(mR_{n})^{n}\,,\\[2.0pt] \text{ if }mR_{n}\in\big[1-\tfrac{1}{n},1+\tfrac{1}{n}\big]&\qquad\mathbf{E}[B_{n}]\asymp n^{-1/s}=n^{-(p-1)}\,,\\[2.0pt] \text{ if }mR_{n}\in\big[1+\tfrac{1}{n},2]&\qquad\mathbf{E}[B_{n}]\asymp((mR_{n})^{s}-1)^{1/s}\,.\end{array}$ (1.12)
@@
Let $s>0$ and assume that $Z\sim\mu$ admits a finite moment of order $q=s+1$. Let $(B_{n}(u))_{u\in T_{n}}$ be defined as in (1.8) and assume that (1.9) holds. Then, if $\limsup_{n\to\infty}mR_{n}\leqslant 1$ and $\inf_{n\geqslant 0}R_{n}>0$, we have, as $n\to\infty$
$\frac{B_{n}}{\mathbf{E}[B_{n}]}\xrightarrow[n\to\infty]{L^{q}}W\,$,
where $W:=\lim_{n\to\infty}\frac{1}{m^{n}}Z_{n}$ is the a.s. and $L^{q}$ limit of the usual martingale associated to the branching process [^3] If $Z\sim\mu$ admits a finite moment of order $q$, the $L^{q}$ convergence of the martingale is non-trivial: we refer to [27, Prop. 1.3] for a proof.. If $R_{n}\equiv R$ with $R\in(0,m^{-1}]$, then the convergence also holds almost surely.
@@
Let $s>0$ and assume that $Z\sim\mu$ admits a finite moment of order $q=s+1$. Let $(B_{n}(u))_{u\in T_{n}}$ be defined as in (1.8) and assume that (1.9) holds and that the following limit exists
$\kappa_{g}:=\lim_{x\downarrow 0}\frac{1}{x^{s+1}}(x-g(x))\in(0,+\infty)\,.$
(Note that for $g(x)=\frac{x}{(1+x^{s})^{1/s}}$, we have $\kappa_{g}=s^{-1}$). Then, if $\lim_{n\to\infty}mR_{n}=\vartheta\in(0,1]$, we have
$\mathbf{E}[B_{n}]\sim\alpha_{s}(\vartheta)\;(a_{n})^{-1/s}\,,\qquad\text{ as }n\to\infty\,$,
for some constant $\alpha_{s}(\vartheta)$, with $\alpha_{s}(1):=(s\kappa_{g}\mathbf{E}[W^{s+1}])^{-1/s}$ in the case $\lim_{n\to\infty}mR_{n}=1$.
Let us mention that Proposition 1.9 improves in particular the result [10, Thm. 1.2], which considers some additional source of randomness on the resistances but treats only the case of linear ($p=q=2$) conductances in the critical case $mR_{n}\equiv 1$ and requires a moment of order $3$, i.e. $q+1$, for the branching process.
1.3.2 Case of an infinite moment of order $q=s+1$
In the case of an infinite moment of order $q$, we need to assume some (either upper or lower) bounds on the truncated $q$-moment of $Z\sim\mu$. More precisely, we assume that there is some $\alpha\in(1,q]$ and some slowly varying function $L(\cdot)$ such that, for $x\geqslant 1$
$\mathbf{E}[(Z\wedge x)^{q}]\geqslant c_{1}L(x)x^{q-\alpha}\,$, (1.13)
$\mathbf{E}[(Z\wedge x)^{q}]\leqslant c_{2}L(x)x^{q-\alpha}\,$, (1.14)
@@ 1.3.2 Case of an infinite moment of order $q=s+1$
Let $h$ be a increasing function such that $h(x)\sim L(1/x)x^{\alpha-1}$ as $x\downarrow 0$, and denote its inverse by $h^{-1}$. Then, $h^{-1}(x)\sim\tilde{L}(x)x^{1/(\alpha-1)}$ as $x\downarrow 0$, for some slowly varying function $\tilde{L}$, see [8, §1.5.7]. In analogy with (1.12), which corresponds to the case $h(x)=x^{q-1}$, define a normalizing sequence $\gamma_{n}$ as follows:
 $\begin{array}[]{lll}\text{ if }mR_{n}\in\big(0,1-\tfrac{1}{n}\big]&\qquad\gamma_{n}=h^{-1}(1-mR_{n})\,(mR_{n})^{n}\,,\\[2.0pt] \text{ if }mR_{n}\in\big[1-\tfrac{1}{n},1+\tfrac{1}{n}\big]&\qquad\gamma_{n}=h^{-1}(1/n)\,,\\[2.0pt] \text{ if }mR_{n}\in\big[1+\tfrac{1}{n},+\infty)&\qquad\gamma_{n}=h^{-1}(mR_{n}-1)\,.\end{array}$ (1.15)
Let us also define $\tilde{\gamma}_{n}$ by setting
 $\begin{array}[]{ll}\text{ if }mR_{n}\in\big(0,1-\tfrac{1}{n}\big]&\qquad\tilde{\gamma}_{n}=h^{-1}(1/n)(n(1-mR_{n}))^{1/s}(mR_{n})^{n}\,,\\[2.0pt] \text{ if }mR_{n}\in\big(1-\tfrac{1}{n},+\infty)&\qquad\tilde{\gamma}_{n}=\gamma_{n}\,.\end{array}$ (1.16)
In practice, in the case $R_{n}\equiv R$ and for $L(x)\equiv 1$ in (1.13)-(1.14), we have:
(i) $\gamma_{n}\asymp(mR)^{n}$ and $\tilde{\gamma}_{n}\asymp n^{\frac{1}{q-1}-\frac{1}{\alpha-1}}\gamma_{n}$ if $mR<1$;
(ii) $\gamma_{n}=\tilde{\gamma}_{n}\asymp n^{-1/(\alpha-1)}$ if $mR=1$;
(iii) $\gamma_{n}=\tilde{\gamma}_{n}\asymp(mR-1)^{1/(\alpha-1)}$ if $mR>1$.
@@
Assume that $0<\inf_{n}mR_{n}\leqslant\sup_{n}mR_{n}\leqslant K$ for some $K<\infty$, and let $\gamma_{n},\tilde{\gamma}_{n}$ be as in (1.15)-(1.16). Then there are constants $c,c^{\prime}$ such that:
@@
Let us also note that the conditions (1.13)-(1.14) are weaker than the tail conditions (1.17)-(1.18). Indeed, on one hand, we have $\mathbf{E}[(Z\wedge x)^{q}]\geqslant x^{q}\mathbf{P}(Z>x)$, so (1.14) implies the upper bound $\mathbf{P}(Z>x)\leqslant c_{2}L(x)x^{-\alpha}$, but this is not optimal in the case $\alpha=q$. However, in the case $\alpha<q$, (1.14) is equivalent to (1.18) with $\hat{L}=L$. On the other hand, to obtain a lower bound on $\mathbf{P}(Z>x)$, one needs to have $\alpha<q$ and both (1.13)-(1.14); simply write $\mathbf{E}[(Z\wedge Ax)^{q}]\leqslant(Ax)^{q}\mathbf{P}(Z>x)+\mathbf{E}[(Z\wedge x)^{q}]$.
@@ 1.4 Organisation of the paper
In section 2 we introduce the main motivation for our results on non-linear conductances: the random cluster model (RCM) on Galton–Watson trees. In Section 2.2, we present the tree recursion arising in the model (we postpone the proof to Appendix A) and, in Section 2.3, we apply our main results on $p$-conductances to get as a corollary the critical behavior of the RCM on a quenched Galton–Watson tree.
@@
  • In Section 6, we prove Theorem 1.10, which treats the moments of $C_{n}$ in the case $Z\sim\mu$ has an infinite moment of order $p$.
  • In Appendix A we prove the RCM tree recursion presented in Section $2.$
  • In Appendix B we provide some technical proofs regarding branching processes with heavy tails, which are necessary for the estimates of Theorem 1.10.
2 The (critical) random cluster model on a quenched Galton Watson tree
The Random Cluster Model (RCM) or FK-percolation, introduced by Fortuin and Kasteleyn in [21], unifies a number of models in statistical physics, including the percolation model, the Ising or $\mathtt{q}$-states Potts models and the uniform spanning tree. Some of these models, when considered on trees, exhibit recursions as in (1.4): this is true in particular for the connection (sometimes called survival) probability of the RCM on a tree, see (2.2). We give a quick presentation of the model below, but we refer to [23] for a more complete overview of the RCM and its connection to other models; see also [20, Ch. 1]. Our goal here is to give the exact asymptotic behavior of the connection probability, both at and near criticality. We obtain in particular that the so-called critical one-arm exponent is either equal to $1$ when $\mathtt{q}\in(0,2)$ (a “percolation regime”) or to $1/2$ when $\mathtt{q}=2$ (for the Ising model, which to our knowledge was studied at criticality only in [26], on $d$-regular trees); note that these exponents agree with the mean-field exponents.
Let us mention that one of the motivation for considering statistical mechanics models on trees is to apply the results to general random tree-like graphs, see e.g. [38] for an overview (see also [14, 18] in the context of the Ising model, and [9, 15, 25] for the random cluster model).